OpenStudy (anonymous):

A triangle with base of 6cm and an area of 15cm^2 has a height of?

OpenStudy (anonymous):

A=1/2 x base x height or A = 1/2 b h Plug in what you know, solve for h.

OpenStudy (anonymous):

I'm guessing 5cm?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

cool